OpenStudy (gabylovesyou):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (gabylovesyou):

y = 1/3x - 8 y = -3x + 12 y = 3x - 24 y = -1/3x - 4

OpenStudy (gabylovesyou):

@chenna I have to go can u give me the answer fast?

OpenStudy (anonymous):

this is a test for u :P

OpenStudy (gabylovesyou):

please.... i begg!!!! :(

OpenStudy (gabylovesyou):

by Private message.... please